# MobilityGPT: Enhanced Human Mobility Modeling with a GPT model

### v2

@@ 4.1 Generative Transformers
By formulating human mobility trajectories as a sequence generation task with road links from $D(V,E)$, transformer-based models can be applied to generate realistic and contextually relevant synthetic trajectories given the real trajectories. The tokenization of road links, combined with the self-attention mechanism, enables the model to capture spatial dependencies and generate coherent and meaningful sequences of road links.
4.2 Tokenizer Modeling
@@
In MobilityGPT, tokens symbolize locations in terms of road links, and the contextual embeddings derived from tokenization contribute to generating embeddings that consider the holistic context of the entire trajectory. For each trajectory, we incorporate an “end of trajectory” token, denoted as <EOT>, which is appended to each trajectory during training. Given a trajectory sequence $\phi\in\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$ in terms of road links, where $\phi_{i}$ denotes the $i$-th token in the trajectory. The tokenization process is defined as follows:
@@
This inclusion of <EOT> serves a dual purpose. First, it acts as a sentinel token to signify the end of a trajectory. This is expressed as if $\text{if }\phi_{i}=\texttt{<EOT>}\text{, then trajectory }i+1\text{ begins}$. The model leverages this information to capture the distinct OD points of each trajectory effectively. Second, this tokenization allows for generating diverse and randomized trajectory sequences. A detailed discussion can be found in Appendix C.3.
4.3 Map Understanding with Gravity Model and Road Connectivity Matrix
Transformer-based models capture sequential information within long data sequences. However, human mobility modeling presents distinct challenges, necessitating a comprehensive understanding of geographical attributes and constraints. For instance, the sequence of links does not contain hidden attributes of human mobility, such as regional mobility flows (traveling from eastbound of city to west), traffic rules (must right-turn, no U-turn), and trip length. We have developed two innovative methods to overcome these challenges: the Gravity Model and the Road Connectivity Matrix (RCM). The Gravity Model enhances the transformer’s capability to understand the mobility flow patterns between regions, providing a context of spatial importance. Together, RCM allows the transformer to learn the complex structure of the road network $D(V,E)$. These strategies collectively train the transformer with an understanding of geospatial information without needing auxiliary data.
$\text{Gravity}(r_{x},r_{y})=\frac{\text{RegionW}(r_{x})\times\text{RegionW}(r_{y})}{d^{2}(r_{x},r_{y})}$
where $d^{2}(r_{x},r_{y})$ is the squared Euclidean distance between regions. Gravity is based on the principle that the interaction between two regions is directly proportional to the product of their respective weights and inversely proportional to the square of the distance between them. During training, as shown in Figure 2, a gravity-aware sampler selects trajectory sequences $\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$ based on $\text{Gravity}(r_{x},r_{y})$, with $\phi_{1}\in r_{x}$ and $\phi_{n}\in r_{y}$.
2) Road connectivity matrix. The clarity of road connectivity is often diminished when using GPS trajectories compared to traditional 2D maps. This is because, in some instances, roads that appear connected on a map might not be so in reality due to physical barriers or design differences, such as one-way streets and pedestrian zones. To mitigate this issue, we propose a method to precisely extract road connectivities from real-world data and incorporate this information into the training process of MobilityGPT. To achieve this, we define the Road Connectivity Matrix (RCM) based on the links from link trajectories, denoted as $\Phi$:
$RC_{\phi_{x},\phi_{y}}=\begin{cases}1&\text{if }(\phi_{x},\phi_{y})\text{ is aconsecutive pair in }\Phi,\\ &\quad\text{or}\\ &\text{if }\phi_{x}\text{ or }\phi_{y}\text{ is in the last row/column,}\\ 0&\text{otherwise}.\end{cases}$ (4)
RCM effectively represents the connectivity between different road links. If two links, $\phi_{x}$ and $\phi_{y}$, are directly connected in the real world, their corresponding element in RCM is set to 1; otherwise, it is set to 0. Additionally, one row and one column, both filled with 1s, are added to the matrix to facilitate the generation of <EOT>. Specifically, the last linear layer of the transformer, denoted as $F$, is adjusted by the RCM prior to the softmax activation function:
@@ 4.3 Map Understanding with Gravity Model and Road Connectivity Matrix
4.4 Reinforcement learning from trajectory feedback (RLTF)
@@
Fine-tuning is essential for sequence generative models (Zhang et al., 2023). Pre-trained transformer architectures provide a foundation, but fine-tuning customizes models to specific objectives, such as similarity of trajectories, enabling them to learn complex patterns and nuances. This process is particularly vital in applications like human mobility, where capturing intricate movement patterns is crucial. Preference datasets are crucial for fine-tuning, allowing domain knowledge and critique to be integrated into training. Building on this, we propose a novel approach to construct a trajectory-aware preference dataset that combines established fine-tuning strategies with a new preference feedback dataset - all without requiring human labeling. This enables the model to self-improve based on its output feedback, minimizing human intervention and addressing scalability and bias challenges (Ouyang et al., 2022).
@@
Reinforcement learning has proven to be a powerful approach for capturing the long-term dependencies and spatiotemporal characteristics inherent in human mobility modeling, owing to the complexity of this task (Reed et al., 2022). The proposed approach, incorporating our novel preference dataset with reinforcement learning, termed Reinforcement Learning from Trajectory Feedback (RLTF), in an automated feedback process, involves several key steps. First, a reward dataset is created by pairing prompts (partial trajectories) with compilations (complete trajectories generated by the model) and assigning scores based on the similarity of the compilation lengths to the original trajectory lengths (see Figure 3). This reward dataset is then utilized to train a reward model $\mathbf{U}$ through supervised learning. Subsequently, the trained reward model $\mathbf{U}$ is employed to fine-tune the MobilityGPT model through a Proximal Policy Optimization (PPO) policy $\pi$, effectively leveraging reinforcement learning without requiring human feedback.
@@
Preference dataset construction for RLTF. Instead of relying on human feedback in fine-tuning of MobilityGPT, we propose using quantitative metrics to construct the preference dataset. Given a partial trajectory prompt, we generate two complete trajectories and assess their similarity to the actual trajectory by assigning a trajectory similarity score based on the relative lengths of the generated and reference trajectories from the training set. This score measures how well the generated trajectories match the characteristics, particularly length, of the training data, serving as a proxy for human preference without requiring human labeling.
@@
The process, as described in Figure 3, begins with a full trajectory $\Phi={\phi_{1},\phi_{2},\ldots,\phi_{n}}$ of length $L(\Phi)$. The first $m$ road-links $\Phi^{\prime}_{0}={\phi_{1},\phi_{2},\ldots,\phi_{m}}$ serve as a prompt for MobilityGPT to predict two possible continuations: $\Phi^{\prime}_{1}={\phi_{m+1},\ldots,\phi_{k}}$ and $\Phi^{\prime}_{2}={\phi_{m+1},\ldots,\phi_{z}}$. These are concatenated with $\Phi^{\prime}_{0}$ to form two complete trajectories, evaluated using $\Gamma_{1}=|L(\Phi)-L(\Phi^{\prime}_{0}+\Phi^{\prime}_{2})|$ and $\Gamma_{2}=|L(\Phi)-L(\Phi^{\prime}0+\Phi^{\prime}1)|$. The trajectory with lower $\Gamma$ is labeled $\Phi_{\text{chosen}}$, and the other $\Phi_{\text{rejected}}$, forming a preference dataset for RLTF. By leveraging this proximity-based metric, we eliminate the need for human labeling, streamlining the dataset creation process and potentially mitigating biases introduced by human subjectivity. The details of reward model training and RL-based fine-tuning methods are explained in Appendix A with more details. It is also worth noting that MobilityGPT can be generalized to different fine-tuning strategies with the proposed preference dataset.
5 Experiments
MobilityGPT’s effectiveness is validated through extensive experiments on real-world datasets against recent benchmarks. We will release our codes and data upon acceptance.
5.1 Experimental setup
1) Dataset description. We have used two datasets of GPS trajectories from taxi drivers in Porto, Portugal, and Beijing, China. The Porto dataset has a sampling period of $15$ seconds, whereas the Beijing dataset has a sampling period of $1$ minute. Both datasets have been preprocessed and map-matched by (Jiang et al., 2023) to remove any errors. The Porto dataset contains $695,085$ unique trips on $11,095$ road links, and the Beijing dataset contains $956,070$ trajectories on $40,306$ road links.
2) Benchmark models. We assess the performance of MobilityGPT with respect to four synthetic trajectory generation models as baselines. We carefully selected the benchmarks as one from each distinct approach: discrete sequential GAN (SeqGAN (Yu et al., 2017)), continuous GPS generation (LSTM-TrajGAN (Rao et al., 2020)), two stage generative model (TS-TrajGen (Jiang et al., 2023)) and continuous diffusion generative model (DiffTraj (Zhu et al., 2024)). The details of these benchmark models can be found in Appendix B.1.
Table 1: Comparison of the utility metrics with benchmark studies. Bold and underlined results show the best and second-best results, respectively. For all metrics, lower values indicate better performance. While query error relies on normalized error rates for most visited places, other metrics use the Jensen-Shannon distributional divergence statistic.
  Porto-Taxi BJ-Taxi
SeqGAN 0.138 0.116 0.220 0.191 0.326 0.063 0.296 0.162 0.275 0.120
DiffTraj 0.267 0.177 0.126 0.123 0.252 0.074 0.146 0.158 0.123 0.114
5.2 Evaluation metrics
Our first utility metric, mainly used for graph data, is query error, which quantifies the mobility characteristics of frequently visited places (Chen et al., 2011). A lower query error indicates that the output synthetic trajectories are more similar to the training trajectories.
The Jensen-Shannon divergence (JSD) is a widely recognized similarity metric primarily employed for quantifying the similarity between two probability distributions (Lin, 1991). Additionally, we evaluate our proposed MobilityGPT using JSD for four metrics: OD similarity, trip length, travel radius, and gravity distribution. Please refer to Appendix B.2 for a more detailed description of our evaluation metrics.
In addition to divergence metrics, we introduce a new metric that measures link trajectory connectivity as a percentage. We employed the connectivity metric in our ablation studies to assess the impact of various components on link trajectory connectivity, as shown in Table 5.4a.
5.3 Comparison with benchmark generative methods
Table 1 summarizes the performance of MobilityGPT and benchmark models. We sampled $5000$ random trajectories from the test set and generated same number of synthetic trajectories. The results show the significant advantage of MobilityGPT over existing state-of-the-art models for trajectory generation. Across both Porto-Taxi and BJ-Taxi, MobilityGPT consistently outperforms all baseline models, achieving the best or second-best performance on multiple utility metrics.
Notably, MobilityGPT demonstrates up to $24.07\%$ improvement over the second-best model in modeling trip lengths and up to $17.07\%$ improvement in capturing spatial radius on BJ-Taxi. Additionally, it exhibits up to $9.27\%$ improvement in capturing gravity patterns on Porto-Taxi. MobilityGPT excels in accurately modeling intricate characteristics of individual trajectories, such as spatial and temporal patterns, by effectively capturing the relationships between locations and trajectories, addressing a significant limitation of previous approaches. Its remarkable performance in modeling OD distributions highlights its capability to capture underlying patterns governing traffic flow and human mobility. Overall, MobilityGPT’s superior performance across multiple metrics highlights its effectiveness as a comprehensive and powerful trajectory generation model, positioning it as a valuable tool for applications in transportation, urban planning, and mobility analysis.
The primary objective of mobility generative models is to produce trajectories that closely resemble the original data by capturing the overall mobility characteristics, rather than generating exact replicas. Figure 4 depicts the spatial densities generated by baseline methods for Porto City. While all the maps demonstrate that the generated trajectories accurately capture the city’s geographic profiles, MobilityGPT has a clear representation of geospatial density in the generated trajectories compared to the original test trajectories. Clearly, MobilityGPT produces trajectories that closely resemble the originals both in urban areas and the outskirts of the city, surpassing other methods significantly.
5.4 Ablation study
As MobilityGPT incorporates three distinct contributions to the standard GPT model for comprehensively capturing human mobility patterns – the gravity model, RCM, and RL fine-tuning (RLTF) – we performed experiments using different variants of these models, evaluated using three comparison metrics as shown in Table 5.4a. On average, with the additional components of MobilityGPT, we observe a $7.4\%$ improvement in trajectory similarities in terms of ODs and a $19.5\%$ improvement over trip length similarities. Furthermore, without the RCM, MobilityGPT fails to preserve connectivity in the generated trajectories. RCM ensures connectivity by conditioning the generation of the next link on connected links, effectively capturing the spatial constraints inherent in human mobility patterns.
We compare different fine-tuning methods with our proposed preference dataset construction approach: supervised fine-tuning (SFT) and direct preference optimization (DPO) (Rafailov et al., 2024) and summarized their performance in Table 5.4b. We found that RLTF is effective for capturing the complexity of mobility modeling and outperforms other fine-tuning strategies.
Table 2: Ablation study shows the influence of (a) removing each component on the performance of MobilityGPT and (b) on different methods for fine-tuning.
       RCM Gravity Model RLTF OD Trip L. Conn.
0.123 0.154 0.87
0.122 0.146 0.83
0.120 0.147 1.0
0.114 0.140 1.0
0.114 0.124 1.0
(a)        SFT DPO RLTF OD Trip Length 0.116 0.171 0.115 0.163 0.114 0.124   (b)
(a) Original Trajectories
(b) MobilityGPT
(c) TS-TrajGen
(d) SeqGAN
(e) LSTM-TrajGAN
(f) DiffTraj
6 Discussion
This paper presents a GPT-based generative method for modeling human mobility characteristics. The proposed mechanism considers a multi-objective learning method with pre-training and fine-tuning stages. MobilityGPT pre-training leverages the understanding of human mobility using the string self-attention structures by conditioning the trajectory sampling with gravity model and sequence sampling with road connectivity matrix. To further enhance the MobilityGPT capabilities for generating similar trajectories to real trajectories, we propose a fine-tuning strategy that does not require a human assessment. The novel fine-tuning method trains a reward using the trajectory trip length and employs the reward model as an import to the RL policy gradient method. The experiments validate the effectiveness of the proposed MobilityGPT approach for generating realistic mobility trajectories from real trajectory samples by comparing with four state-of-the-art models.
Limitation: MobilityGPT achieves state-of-the-art results while handling road network constraints, but lacks formal privacy guarantees for synthetic trajectories. Future directions involve improving generalizability and incorporating practical privacy assurances.
Appendix A Reinforcement Learning from Trajectory Feedback Details
This section provides details about the novel RLTF fine-tuning strategy for generating high-quality synthetic mobility trajectories.
A.1 Reward model training.
@@ Appendix A Reinforcement Learning from Trajectory Feedback Details
A Reward (preference) Model $\mathbf{U}$ is trained using the proposed preference dataset to predict the probability that a given trajectory is chosen over a rejected alternative. The loss function is formulated as follows:
$-\log(\sigma(\mathbf{U}(\Phi_{\text{chosen}})-\mathbf{U}(\Phi_{\text{rejected}})))$ (7)
@@
A.2 Fine-tuning with reinforcement learning.
Given the pre-trained MobilityGPT model and supervised trained reward model, we train an RL policy that optimizes the MobilityGPT model for generating higher-quality synthetic trajectories. Since policy learning relies on the reward that policy receives from the environment, the reward model provides feedback for RL policy optimization. We followed a general policy optimization method as in (Ziegler et al., 2019). The RL agent learns a fine-tuning policy $\pi$, which is optimized with the well-known Proximal Policy Optimization (PPO) method (Schulman et al., 2017). The reward of policy $\pi$ consists of two terms: reward model logits given the prompt $\Phi^{\prime}_{p}$ and compilation $\Phi^{\prime}_{c}$ pair received from the MobilityGPT as $\mathbf{U}(\Phi^{\prime}_{p},\Phi^{\prime}_{c})$ and policy shift constraint in terms of the KL divergence given the baseline policy $\pi$. The full reward can be represented as follows:
@@
Appendix B Experimental Setup Details
This section describes the benchmark models and evaluation metrics in detail.
B.1 Benchmark models

### v3

@@ IV-A Generative Transformers
By formulating human mobility trajectories as a sequence generation task with road links from $D(V,E)$, transformer-based models can be applied to generate realistic and contextually relevant synthetic trajectories given the real trajectories. The tokenization of road links, combined with the self-attention mechanism, enables the model to capture spatial dependencies and generate coherent and meaningful sequences of road links. Next, we explain our unique tokenizer modeling.
IV-B Tokenizer Modeling
Tokenization, the breakdown of sequences into smaller units or tokens, is a key step in preparing raw data for transformer models. The influence of tokenization extends to trajectory sequences, where tokens represent locations or spatiotemporal points.
@@ IV-B Tokenizer Modeling
In MobilityGPT, tokens symbolize locations in terms of road links, and the contextual embeddings derived from tokenization contribute to generating embeddings that consider the holistic context of the entire trajectory.
To structure the input data into a format suitable for autoregressive modeling, each trajectory is treated as a discrete sequence of location-based tokens. For each trajectory, we incorporate an “end of trajectory” token, denoted as <EOT>, which is appended to each trajectory during training. Given a trajectory sequence $\phi\in\Phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$ in terms of road links, where $\phi_{i}$ denotes the $i$-th token in the trajectory. The tokenization process is defined as follows:
$\mathbf{\Phi}_{\text{tokenized}}=\{\phi_{1},\phi_{2},...,\phi_{n},\texttt{<EOT>}\}$ (3)
This inclusion of <EOT> serves a dual purpose. First, it acts as a sentinel token to signify the termination of a trajectory. Formally, $\text{if }\phi_{i}=\texttt{<EOT>}\text{, then trajectory }i+1\text{ begins}$. This structural sequence tokenization enables the model to distinguish between successive locations, thus avoiding positional leakage across trajectories during batch processing. This prevents the model from interpreting the start of a new trajectory in a batch as a continuation of the previous one, maintaining clear trajectory boundaries during both training and generation. Second, the proposed tokenization supports the generation of diverse and randomized trajectory sequences by clearly marking sequence endpoints, which facilitates autoregressive sampling without conflating distinct OD pairs. As <EOT> is part of the model’s vocabulary, MobilityGPT assigns it a probability at every decoding step. This allows trajectories to terminate at variable positions, enabling the generation of diverse and randomized geospatial sequences while preserving structural coherence with real-world mobility patterns.
Moreover, this design abstracts away from raw GPS coordinates, allowing the model to learn high-level movement patterns through a discrete vocabulary of meaningful road link identifiers. This abstraction mitigates the impact of noise from GPS sparsity or irregular sampling rates and grounds the representation in topological connectivity rather than continuous space.
IV-C Map Understanding with Gravity Model and Road Connectivity Matrix
Transformer-based models capture sequential information within long data sequences. However, human mobility modeling presents distinct challenges, necessitating a comprehensive understanding of geographical attributes and constraints. For instance, the sequence of links does not contain hidden attributes of human mobility, such as regional mobility flows (traveling from eastbound of city to west), traffic rules (must right-turn, no U-turn), and trip length. To address these limitations, we integrate two structured priors from classical mobility modeling—the gravity model and the RCM—into the training process of the transformer.
While both models are well-established in the transportation literature [cabanas2025human, wang2023attention], our work recontextualizes them as inductive biases within a deep generative model. Specifically, the gravity model is repurposed as a data sampling prior to expose the transformer to representative and balanced regional flows, while the RCM is injected into the final decision layer of the model to constrain predictions to physically feasible road transitions. These strategies collectively guide the transformer to learn geospatial structure without requiring explicit auxiliary data such as maps or network graphs.
$\text{Gravity}(r_{x},r_{y})=\frac{\text{RegionW}(r_{x})\cdot\text{RegionW}(r_{y})}{d^{2}(r_{x},r_{y})}$ (4)
where $d^{2}(r_{x},r_{y})$ is the squared Euclidean distance between regions. This classical formulation has historically been used to model inter-regional travel demand. In our case, we adopt it as a training-time sampling prior to ensure that the model observes a diverse and flow-consistent distribution of OD pairs. During training, as shown in Figure 2, a gravity-aware sampler selects trajectory sequences $\Phi={\phi_{1},\phi_{2},...,\phi_{n}}$ based on $\text{Gravity}(r_{x},r_{y})$, where $\phi_{1}\in r_{x}$ and $\phi_{n}\in r_{y}$. This strategy reduces sampling bias toward overrepresented regions and encourages the model to internalize global mobility trends.
2) Road connectivity matrix. The clarity of road connectivity is often diminished when using GPS trajectories compared to traditional 2D maps. This is because, in some instances, roads that appear connected on a map might not be so in reality due to physical barriers or design differences, such as one-way streets and pedestrian zones. To incorporate real-world structural constraints into sequence generation, we propose the RCM, a sparse binary matrix that reflects the physical adjacency of road links.
Let $\Phi$ denote the set of training trajectories in terms of road links. The RCM is defined as:
@@ IV-C Map Understanding with Gravity Model and Road Connectivity Matrix
RCM effectively represents the connectivity between different road links. If two links, $\phi_{x}$ and $\phi_{y}$, are directly connected in the real world, their corresponding element in RCM is set to 1; otherwise, it is set to 0. This structure reflects only plausible transitions observed in the data. To enforce these structural constraints consistently, we apply the RCM to the transformer’s output logits during both training and generation, ensuring that the model only learns and predicts physically plausible road transitions. Additionally, one row and one column, both filled with 1s, are added to the matrix to facilitate the generation of <EOT>. Specifically, the last linear layer of the transformer, denoted as $F$, is adjusted by the RCM prior to the softmax activation function:
@@
IV-D Reinforcement learning from trajectory feedback (RLTF)
@@
Fine-tuning is essential for sequence generative models [zhang2023instruction]. Pre-trained transformer architectures provide a foundation, but fine-tuning customizes models to specific objectives, such as similarity of trajectories, enabling them to learn complex patterns and nuances. This process is particularly vital in applications like human mobility, where capturing intricate movement patterns is crucial. Building on this, we propose a novel approach to construct a trajectory-aware preference dataset that combines established fine-tuning strategies with a new preference feedback dataset - all without requiring human labeling. This enables the model to self-improve based on its output feedback, minimizing human intervention and addressing scalability and bias challenges [ouyang2022training].
@@
Reinforcement learning has proven to be a powerful approach for capturing the long-term dependencies and spatiotemporal characteristics inherent in human mobility modeling, owing to the complexity of this task [reed2022generalist]. The proposed approach, incorporating our novel preference dataset with reinforcement learning, termed Reinforcement Learning from Trajectory Feedback (RLTF), in an automated feedback process, involves several key steps. First, a reward dataset is created by pairing prompts (partial trajectories) with compilations (complete trajectories generated by the model) and assigning scores based on the similarity of the compilation lengths to the original trajectory lengths (see Figure 3). This reward dataset is then utilized to train a reward model $\mathbf{U}$ through supervised learning. Subsequently, the trained reward model $\mathbf{U}$ is employed to fine-tune the MobilityGPT model through a Proximal Policy Optimization (PPO) policy $\pi$, effectively leveraging reinforcement learning without requiring human feedback.
The reward model $\mathbf{U}$ is a crucial component of our RLTF framework. We adopted the same lightweight and efficient model architecture as MobilityGPT itself for the reward model. This design choice ensures computational efficiency and a seamless integration with the main model, allowing the reward signal to be accurately modeled without introducing significant overhead. By using a shared architecture, we can effectively capture the complexities of the reward function while keeping the overall system resource-friendly.
@@
Preference dataset construction for RLTF. Instead of relying on human feedback in fine-tuning of MobilityGPT, we propose using quantitative metrics to construct the preference dataset. Given a partial trajectory prompt, we generate two complete trajectories and assess their similarity to the actual trajectory by assigning a trajectory similarity score based on the relative lengths of the generated and reference trajectories from the training set (See Step 2 in Figure 2). This score measures how well the generated trajectories match the characteristics, particularly length, of the training data, serving as a proxy for human preference without requiring human labeling.
@@
The process, as described in Figure 3, begins with a full trajectory $\Phi={\phi_{1},\phi_{2},\ldots,\phi_{n}}$ of length $L(\Phi)$. The first $m$ road-links $\Phi^{\prime}_{0}={\phi_{1},\phi_{2},\ldots,\phi_{m}}$ serve as a prompt for MobilityGPT to predict two possible continuations: $\Phi^{\prime}_{1}={\phi_{m+1},\ldots,\phi_{k}}$ and $\Phi^{\prime}_{2}={\phi_{m+1},\ldots,\phi_{z}}$. These are concatenated with $\Phi^{\prime}_{0}$ to form two complete trajectories, evaluated using $\Gamma_{1}=|L(\Phi)-L(\Phi^{\prime}_{0}+\Phi^{\prime}_{2})|$ and $\Gamma_{2}=|L(\Phi)-L(\Phi^{\prime}0+\Phi^{\prime}1)|$. The trajectory with lower $\Gamma$ is labeled $\Phi_{\text{chosen}}$, and the other $\Phi_{\text{rejected}}$, forming a preference dataset for RLTF. By leveraging this proximity-based metric, we eliminate the need for human labeling, streamlining the dataset creation process and potentially mitigating biases introduced by human subjectivity. MobilityGPT can be generalized to different fine-tuning strategies with the proposed preference dataset. In Figure 4, we exemplified one step of the preference dataset creation schema with an actual sample trajectory using the Porto dataset [jiang2023continuous].
Figure 4: Example of a preference dataset, showing a sample input alongside two completed trajectory pairs. Candidate 1’s trajectory demonstrates greater similarity to the Baseline trajectory compared to Candidate 2.
@@
Reward model training. A Reward (preference) Model $\mathbf{U}$ is trained using the proposed preference dataset to predict the probability that a given trajectory is chosen over a rejected alternative (See Step 2 in Figure 2). The loss function is formulated as follows:
$\log(\sigma(\mathbf{U}(\Phi_{\text{chosen}})-\mathbf{U}(\Phi_{\text{rejected}})))$ (8)
@@
Fine-tuning with reinforcement learning. Given the pre-trained MobilityGPT model (Step 1) and supervised trained reward model (Step 2), we train an RL policy that optimizes the MobilityGPT model for generating higher-quality synthetic trajectories ss depicted in Step 3 of Figure 2. Since policy learning relies on the reward that policy receives from the environment, the reward model provides feedback for RL policy optimization. We followed a general policy optimization method as in [ziegler2019fine]. The RL agent learns a fine-tuning policy $\pi$, which is optimized with the well-known Proximal Policy Optimization (PPO) method [schulman2017proximal]. The reward of policy $\pi$ consists of two terms: reward model logits given the prompt $\Phi^{\prime}_{p}$ and compilation $\Phi^{\prime}_{c}$ pair received from the MobilityGPT as $\mathbf{U}(\Phi^{\prime}_{p},\Phi^{\prime}_{c})$ and policy shift constraint in terms of the (Kullback–Leibler) KL divergence given the baseline policy $\pi$. The full reward can be represented as follows:
@@
V Experiments
MobilityGPT’s effectiveness is validated through extensive experiments on real-world datasets against recent benchmarks. We performed all the experiments on a single NVIDIA TITAN RTX GPU with 24GB memory. The source code of MobilityGPT is publicly available. The preprint version of this paper is available on arXiv at [haydari2024mobilitygpt].  [^1] https://github.com/ammarhydr/MobilityGPT
V-A Experimental setup
We have used two datasets of GPS trajectories from taxi drivers in Porto, Portugal, and Beijing, China. The Porto dataset has a sampling period of $15$ seconds, whereas the Beijing dataset has a sampling period of $1$ minute. Both datasets were preprocessed and map-matched following the protocol described by [jiang2023continuous]. Specifically, the authors filtered out abnormal trajectories by removing those shorter than five road links or containing loops. After preprocessing, the Porto dataset comprises $695,085$ unique trips across $11,095$ road links, while the Beijing dataset includes $956,070$ trajectories spanning $40,306$ road links.
We chose these two datasets for two primary reasons. First, they are widely recognized and frequently used as benchmark datasets in the field of human mobility modeling, particularly in recent works such as TS-TrajGen [jiang2023continuous] and DiffTraj [zhu2024difftraj]. This established use allows for a direct and fair comparison of our model’s performance against state-of-the-art benchmarks. Second, the datasets represent distinct geographic regions and mobility patterns. By validating our model on data from different continents with varying urban structures, we aim to demonstrate its robustness and ability to accurately model diverse human mobility characteristics, irrespective of geographical location or a specific urban context.
V-B Benchmark models